### v1

@@ 3 Results
Lastly, we perform the same analysis on the the least ($L_{z}=13.33h$) and the most ($L_{z}=0.83$) constrained jets at several locations on the jet centerline ($x=40h,80h,120h$) (see Fig. 3b, c). In these cases, $S_{3}$ has the same sign throughout $x$, indicating that turbulence is either three- ($S_{3}<0$) or two- ($S_{3}>0$) dimensional in the planar jet. The structure functions show power-law scaling for intermediate values of the separation distance $r$ in the three-dimensional planar jet; the scaling law $S_{3}\sim-r$ agrees with the direct cascade of energy. On the other hand, the scaling for $S_{3}$ in the two-dimensional planar jet shifts towards $S_{3}\sim r^{3}$, indicating instead the presence of the direct cascade of enstrophy. More interestingly, the scaling holds for small separation distances farther from the inlet. This is in good agreement with the concept of the enstrophy cascade as a space-filling phenomenon, thus being present at very small scales (Benzi et al., 1986). Additionally, $S_{3}$ follows an anomalous behavior at large and intermediate $r$. As observed in Fig. 1, large-size coherent vortices emerge far away from the inlet, which interrupt the enstrophy cascade. At these distances, turbulence becomes more intermittent as velocity fluctuations are localized in the vortexes.
@@ 4 Conclusions
We have studied via direct numerical simulations how the vertical confinement of a turbulent planar jet can alter its dimensionality. We show that, under the right constraint, a mixed-dimensional turbulent regime appears, that is characterized by the simultaneous presence of large-size two-dimensional and small-size three-dimensional scales. The onset of this particular regime is dictated by the size of the constrained dimension $L_{z}$: as soon as the flow scales become larger than $L_{z}$, the flow dimensionality changes from three- to two- dimensions. This transition is postponed further downstream as the flow is less constrained (increasing $L_{z}$). Therefore, for sufficiently large $L_{z}$, turbulent scales are simply three-dimensional along the jet, and the direct cascade of energy is enabled. Conversely, a strong confinement (small $L_{z}$) makes the flow two-dimensional: the direct cascade of energy is disrupted and the direct cascade of enstrophy takes place. Both cascades are conserved wherever the mixed-dimensional turbulent state is present: the direct cascade of energy is active at small scales, whereas the direct cascade of enstrophy dominates at large scales. In addition, we found that the mixed-dimensional configuration is an overall more energetic state, thus partially retaining the three-dimensionality in the flow while deferring the emergence of two-dimensional strong vortical structures further downstream from the inlet. The direct enstrophy and energy cascades and the respective scalings we report here are the same of a Newtonian fluid: the Carreau fluid is a non-Newtonian fluid model characterized by shear-thinning alone, which allows to attain Newtonian turbulence at a relatively low Reynolds number.
@@
[Acknowledgements]The research was supported by the Okinawa Institute of Science and Technology Graduate University (OIST) with subsidy funding from the Cabinet Office, Government of Japan. The authors acknowledge the computational resources provided by the Scientific Computing section of the Research Support Division at OIST.

### v4

@@ 4.1 Effect of vertical confinement
Lastly, we perform the same analysis on the least ($L_{z}=13.33h$) and the most ($L_{z}=0.83$) constrained jets at several locations on the jet centerline ($x=40h,80h,120h$) (see Fig. 3b, c). In these cases, $S_{3}$ has the same sign throughout $x$, indicating that turbulence is either three- ($S_{3}<0$) or two- ($S_{3}>0$) dimensional in the planar jet. For the three-dimensional planar jet, the structure function approaches the scaling for the direct cascade of energy, $S_{3}\sim-r$, at intermediate values of the separation distance $r$. On the other hand, the scaling for $S_{3}$ in the two-dimensional planar jet shifts towards $S_{3}\sim r^{3}$, indicating instead the presence of the direct cascade of enstrophy. More interestingly, the scaling holds for small separation distances further from the inlet. This is in good agreement with the concept of the enstrophy cascade as a space-filling phenomenon, thus being present at very small scales (Benzi et al., 1986). Additionally, $S_{3}$ follows an anomalous behavior at large and intermediate $r$. As observed in Fig. 1, large-size coherent vortices emerge far away from the inlet, which interrupt the enstrophy cascade. At these distances, turbulence becomes more intermittent as velocity fluctuations are localized in the vortexes.
4.2 Effect of $Re$ and $Cu$ in $2.5$D turbulence
a)b)c)d)e)
Figure 4: Effect of $Re$ and $Cu$ in the mixed-dimensional turbulent planar jet ($L_{z}=3.33h$). Panels a and b report the power spectrum computed at the jet centerline at distances $x=40h,80h,120h$ from the inlet. The 3D turbulence scaling $f^{-5/3}$ (red) and the 2D turbulence scaling $f^{-3}$ (blue) are reported for reference. Panels c and d: third-order structure function $S_{3}$ at increasing stream-wise distances, $x=40h,80h,120h$. Red markers identify three-dimensional turbulence and blue markers two-dimensional turbulence. The expected scalings, $r$ (3D) and $r^{3}$ (2D) are reported for reference. Panel e: transition scale $r$ at which $S_{3}$ changes sign as a function of the stream-wise position. Markers identify the different cases: squares for $Re=40$, diamonds for $Cu=200$ and circles for the reference case ($Re=20$, $Cu=100$). The color of the marker defines the dimensionality of the flow: red for $3$D ($S_{3}<0$), blue for $2$D ($S_{3}>0$) and purple for $2.5$D ($S_{3}$ changes sign). The inset reports the transition scale $r$ normalized by the local jet thickness, $2\delta$.
In the previous section we observed that constraining the flow in the vertical direction leads to a transition in the turbulent flow, from three-dimensional turbulence (large $L_{z}$), to a mixed $2.5$D state (intermediate $L_{z}$) and lastly to two-dimensional turbulence (low $L_{z}$). We now consider an intermediate case, $L_{z}=3.33h$, characterized by mixed-dimensional (2.5D) turbulence and investigate the sensitivity of the flow dimensionality on the problem parameters. We separately test the effect of a flow parameter, the inlet Reynolds number $Re$, and of the fluid rheology. Specifically, we consider a case at a higher Reynolds number, $Re=40$, and a case at a higher Carreau number, $Cu=200$. An increase in either parameter results in a higher local Reynolds number. The increase in the inlet Reynolds number is attained by halving the reference, zero-shear viscosity $\mu_{0}$ (the infinite-shear viscosity is reduced accordingly, $\mu_{\infty}=\mu_{0}/50$), whereas the Carreau number is increased by doubling the fluid consistency index $\lambda$ with all other parameters left unchanged. The effect of the power-law index $n$ was not tested, as a reduction of $n$ leads to negligible changes in the fluid rheology and an increase of $n$ leads to a laminarization of the jet fluid, due to an increase in the local viscosity. An increase in the Carreau number shifts the transition towards the infinite-shear viscosity at lower values of the local shear rate $\dot{\gamma}$.
We report data from these two additional cases in Fig. 4, with $Re=40$ in the left column and $Cu=200$ in the right column. The turbulent kinetic energy power spectrum, panels a and b, shows similar results to the reference case ($L_{z}=3.33h$, $Re=20$ and $Cu=100$): three- and two-dimensional regimes coexist (large 2D scales and small 3D scales). Near the inlet, $x=40h$, we observe three-dimensional turbulence characterized by the $f^{-5/3}$ scaling of the energy spectrum; as we move downstream, the vertical constraint forces turbulence to the two dimensional regime, as indicated by the scaling $f^{-3}$. For the case at larger Carreau number, $Cu=200$, we however observe a steepening of the power spectrum away from the inlet: at $x=120h$, the power spectrum approaches the scaling $f^{-5}$, similarly to what observed in the more constrained cases. The cases at either higher $Re$ or $Cu$ are characterized by a widening of the region where the power-law scalings are observed: the wider separation of scales is owed to the higher value of the local Reynolds number. Also the third-order structure functions $S_{3}$, reported in panels c and d at three different stream-wise positions, show that the flow is characterised by a mixed-dimensional regime, similarly to what observed for the reference case, Fig. 3a. We observe that for the $Cu=200$ case the onset of the mixed-dimensional regime occurs at a larger distance from the inlet: at $x=40h$ the structure function is negative at each scale, thus denoting the absence of $2$D turbulence. At further downstream positions, both the $Re=40$ and $Cu=200$ cases show a transition between $2$D and $3$D very similar to the reference case.
The scale at which we observe transition between two- and three-dimensional turbulence is reported in Fig. 4e for different stream-wise positions. This transition scale is defined as the scale $r$ where the third-order structure function changes sign, from $S_{3}<0$ (3D) to $S_{3}>0$ (2D), and corresponds to the largest $3$D scale attained at the specific stream-wise position. We use markers to differentiate the various cases (squares for $Re=40$, diamonds for $Cu=200$ and circles for the reference case) and colors to identify the regime: red for $3$D turbulence ($r$ corresponds to the largest scale, $2\delta$), purple for $2.5$D ($r$ corresponds to the transition scale) and blue for $2$D ($r$ corresponds to the smallest scale considered). As seen in panel d, the case at higher $Cu$ shows a later transition, at $x>60h$, whereas the case at higher $Re$ exhibits very similar transition length-scale to the reference case. The transition from 3D to 2.5D is initiated at a scale similar to twice the jet thickness, $r\approx 1.5\delta$, as shown in the inset in Fig. 4e, and as we move downstream it becomes much smaller than the jet thickness. For the reference and $Re=40$ cases we observe a non-monotonic trend in the transition scale $r$ (purple markers), initially increasing up to $x=60h$ and decreasing past this stream-wise position. The decreasing pattern for $x\geq 70h$ is shown as well by the $Cu=200$ case, which instead shows only the monotonically decreasing trend (for $x\geq 70h$, the first stream-wise position at which mixed-dimensional turbulence is reported). The maximum transition scale, achieved at about $x=70h$, is about $3L_{z}$ for all cases. We report as well a difference in the transition from $2.5$D to $2$D among the three cases: the flow becomes two-dimensional for both the $Cu=200$ (at $x=140h$) and the reference case (at $x=130h$). The $Re=40$ remains instead characterized by mixed-dimensional turbulence and does not show a transition to $2$D turbulence within the length of the simulation domain. Nevertheless, we observe that the size of the transition scale $r$ decays with the stream-wise position, thus indicating that eventually the flow will become two-dimensional at a large enough distance from the inlet. We attribute the difference in the $2.5$D to $2$D transition to the lower viscosity of the fluid: the $Re=40$ is characterized by a lower reference zero-shear (and as well, infinite-shear) viscosity, thus delaying the transition to $2$D turbulence.
@@ 5 Conclusions
We have studied via direct numerical simulations how the vertical confinement of a turbulent planar jet can alter its dimensionality. We show that under the right constraint, a mixed-dimensional turbulent regime appears, that is characterized by the simultaneous presence of large-size two-dimensional and small-size three-dimensional scales. The onset of this particular regime is dictated by the size of the constrained dimension $L_{z}$: as soon as the flow scales become larger than $L_{z}$, two-dimensional flow structures appear. This transition is postponed further downstream as the flow is less constrained (increasing $L_{z}$). Therefore, for sufficiently large $L_{z}$, turbulent scales are simply three-dimensional along the jet, and the direct cascade of energy is enabled. Conversely, a strong confinement (small $L_{z}$) makes the flow two-dimensional: the direct cascade of energy is disrupted and the direct cascade of enstrophy takes place. Both cascades are conserved wherever the mixed-dimensional turbulent state is present: the direct cascade of energy is active at small scales, whereas the direct cascade of enstrophy dominates at large scales. The generality of these findings has been tested: two additional cases, one at a higher inlet Reynolds number and one at a higher Carreau number, have been investigated for the mixed-dimensional case, $L_{z}=3.33$. We report that the mixed-dimensional 2.5D turbulent regime is still observed at both higher $Re$ and higher $Cu$: the flow is still characterized by large two-dimensional and small three-dimensional scales. The largest transition scale, marking the largest 3D turbulent scales, is attained at $x=70h$ for the three cases ($Re=40$, $Cu=200$ and reference case). Minor differences are observed in the transition from 3D to 2.5D, with the $Cu=200$ case showing 2.5D turbulence further downstream, and in the transition from 2.5D to 2D, which does not occur within the computational domain for the $Re=40$ case. We expect that considering a turbulent Newtonian jet (at a comparable inlet local Reynolds number) would have similar effects of increasing the inlet Reynolds number (case at $Re=40$, see figure 4e): the region characterized by $2.5$D turbulence widens and the transition to $2$D flow shifts downstream. A Newtonian fluid lacks any shear-thinning properties, thus it is not characterized by a local viscosity increasing with the stream-wise position, thus allowing smaller turbulent structures to exist over a longer distance from the inlet. To better quantify this effect, a complete comparison with Newtonian turbulence is required, and may be the object of a future study. In addition, when increasing either the Reynolds or the Carreau number, we found that the mixed-dimensional configuration is an overall more energetic state, thus partially retaining the three-dimensionality in the flow while deferring the emergence of two-dimensional strong vortical structures further downstream from the inlet. The direct enstrophy and energy cascades and the respective scalings we report here are the same of a Newtonian fluid: the Carreau fluid is a non-Newtonian fluid model characterized by shear-thinning alone, which allows to attain Newtonian turbulence at a relatively low Reynolds number.
@@
[Acknowledgements]The research was supported by the Okinawa Institute of Science and Technology Graduate University (OIST) with subsidy funding from the Cabinet Office, Government of Japan. M.E.R. also acknowledges funding from the Japan Society for the Promotion of Science (JSPS), grant 24K17210. The authors acknowledge the computer time provided by the Scientific Computing and Data Analysis section of the Core Facilities at OIST, and the computational resources on SQUID provided by the Cybermedia Center at Osaka University through the HPCI System Research Project (project ID: hp230018).
@@
Appendix A Structure function in planar jets
The $p$-th order structure function is defined as the $p$-th moment of velocity differences (Frisch, 1995):
$S_{p}(r)=\langle\left(\Delta u\right)^{p}\rangle.$ (5)
Angle brackets indicate averaging in time, in the vertical direction $z$ and over velocity differences separated by the same distance $r=\left|\boldsymbol{r}\right|$.
The structure function is computed within a cylinder (in red in the sketch in Fig. 5) at a set distance $D$ from the inlet. The cylinder axis is aligned with the vertical direction $z$ and its radius is equal to the jet thickness at distance $D$ from the inlet, i.e. $\delta(D)$. The jet thickness $\delta$ is defined as the distance from the centerline at which the streamwise velocity equals half of the centerline velocity. Longitudinal velocity differences,
$\Delta u(r)=u_{||}\left(\boldsymbol{x}+\boldsymbol{r}\right)-u_{||}\left(\boldsymbol{x}\right)=\left(\boldsymbol{u}\left(\boldsymbol{x}+\boldsymbol{r}\right)-\boldsymbol{u}\left(\boldsymbol{x}\right)\right)\cdot\boldsymbol{r}/\left|\boldsymbol{r}\right|,$ (6)
are computed across points located at distance $r$ lying within the cylinder; structure function data are then averaged over time and over couple of points separated by the same separation distance $r$ (angle brackets in Eq. 5). As the slenderness of the cylinder changes among different vertical constraints and different stream-wise positions (for instance, slender cylinder for $x=40h$ and $L_{z}=13.33h$, or stocky cylinder for $x=120h$ and $L_{z}=0.83h$), we investigated the effect of different cylinder aspect ratios on the computed structure function. The third-order structure function was thus recomputed considering only couple of points lying on the same $x-y$ plane, thus eliminating the effect of the separation in the $z$ direction. Direct comparison among the structure function computed within a cylinder and computed on $x-y$ planes showed minimal differences in the values of the third-order structure function and no difference in its sign.
The jet thickness $\delta$ increases with the distance from the inlet $D$: as we move away from the inlet, the maximum separation distance, $2\delta$, increases as well. The maximum separation distance is divided in $N_{r}=60$ uniformly spaced bins; the width of the bins is thus proportional to the jet thickness, and generally increases with the distance from the inlet $D$. The smallest width of the bin, found close to the inlet, is about the same size of the grid spacing; as one moves away from the inlet the width of the bin increases, resulting in a limited loss of spatial resolution at the smallest scales (nonetheless smaller than ten grid spacings at the farthest distance considered in the present work). Therefore, in some cases very small scales may still show three-dimensional turbulence (and a direct energy cascade), however they may not be fully detected by the structure function, as they are averaged over the smallest separation distance considered.
Figure 5: Calculation of the differences of velocity $\Delta u$ in the planar jet. The differences of longitudinal velocity $u_{||}$ between points separated by a distance $r$ is computed among points within the red cylinder.
In this study, we utilize the third-order structure function $S_{3}$, which we show to be a reliable tool to inspect the flow dimensionality and the preferred turbulent cascade process scale-by-scale. The first derivation of $S_{3}$ resulted in the celebrated Kolmogorov’s 4/5th law in three-dimensional turbulence (Kolmogorov, 1991): $S_{3}(r)=-\frac{4}{5}\epsilon r$, where $\epsilon$ is the mean energy dissipation per unit mass. The $S_{3}$ laws in two-dimensional turbulence were derived in the late 90’s (Lindborg, 1999; Bernard, 1999): for the inverse energy cascade, $S_{3}(r)=\frac{3}{2}Pr$ (where $P$ is the mean energy injection per unit mass), and for the direct enstrophy cascade, $S_{3}(r)=\frac{1}{8}\zeta r^{3}$ (where $\zeta$ is the mean enstrophy dissipation per unit mass). Cerbus & Chakraborty (2017) related $S_{3}(r)$ to the flux functions:
$S_{3}(r)=-\frac{3}{2}\Pi\left(\frac{a}{r}\right)r+\frac{1}{8}Z\left(\frac{a}{r}\right)r^{3}+\cdots,$ (7)
with $\Pi$ and $Z$ being the fluxes of energy and enstrophy, respectively, and $a$ an $O(1)$ numerical constant (the quantity $a/r$ can be thought of as a wavenumber, $k\sim a/r$). Equation 7 expresses $S_{3}(r)$ as a combination of the energy and enstrophy fluxes and its sign is affected by values of both fluxes. The previous laws for $2$D turbulence can be easily derived from Eq. 7 (Cerbus & Chakraborty, 2017).